OpenStudy (anonymous):

Find the domain and range of: f(x)=18sqrt(x)-17

OpenStudy (anonymous):

D={x: where x belong to R and x >=17} R={ y>=0}

OpenStudy (anonymous):

so, would it be 18x-17>=0? and solve?

OpenStudy (anonymous):

is this your function: \(\large f(x)=18\sqrt{x}-17 \)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):



OpenStudy (anonymous):

the domain of f is all the x values that give f(x) a real value.... so looking at the function, the function will not give you a real number when x is less than 0....

OpenStudy (anonymous):

so the actual question in finding the domain is "what x values is \(\large \sqrt x =real\) number

OpenStudy (anonymous):

would I do.. 18x-17<0 and solve?

OpenStudy (anonymous):

x>=-17

OpenStudy (anonymous):

root of x is always larger than 0

OpenStudy (anonymous):

You're just looking to be sure that what's under the root does not go negative. So x cannot be negative. Since the smallest value of x is 0 (because of the domain), then the range's minimum value will happen when x is 0. (So plug in 0 for x, and see what y is). The upper limit of the function will go on forever.

OpenStudy (anonymous):

-17?

OpenStudy (anonymous):

You got it, that's the minimum of the range. As x gets bigger and bigger, there is no limit, it just grows and grows forever.

OpenStudy (anonymous):

so.. [-17,infinity) ?

OpenStudy (anonymous):

That is your range, yes.

OpenStudy (anonymous):

how do I find the domain with this function?

OpenStudy (anonymous):

(-inf,inf) ?

OpenStudy (anonymous):

Well, as we talked about earlier, the domain would be anything that is a "no no". You have a square root, and here, you cannot take the square root of a negative number, as it has no real solutions.

OpenStudy (anonymous):

anything that is NOT a "no-no", sorry ;]

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

No, you simply have sqrt(x), since you cannot take the square root of a negative number, we know that x MUST be greater than or equal to 0.

OpenStudy (anonymous):

so.. 17?

OpenStudy (anonymous):